OpenStudy (anonymous):

Af/ax. f(x,y)=( xy-1)^2

OpenStudy (sanra123):

can u plz help mw with my question

terenzreignz (terenzreignz):

af/ax? Surely you mean \[\Large \frac{\partial f}{\partial x}\] right? :D

OpenStudy (anonymous):

Yes ;)

terenzreignz (terenzreignz):

Well then, it's actually pretty easy. Just differentiate \[\Large (xy^{-1})^2 = x^2y^{-2}\]with respect to x, and pretend y is just a constant.

OpenStudy (anonymous):

=2x(xy-1) ?